News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 14 July 2000 3.30pm EDT

During the last week, Chandra completed the observing schedule as planned, except that the last 15% of NGC 1333 and the second part of two observations of M33 were not done, due to a large solar proton event. The SCS 107 was executed at 11:29am UT on July 13, to protect the science instruments. The team rebuilt and reviewed loads so that the observing program was resumed at 04:47 UT on July 14, in time to perform the first series of observations on Comet c/LINEAR.

Because the normal observing program had been stopped, the ACIS team took the opportunity to execute the 7th in the series of tests of the squeegy mode on July 13. This test explored minimization of the dead time needed for squeegy mode. It had been scheduled to be performed on July 14, which time will now be used to obtain additional data on what looks like the most promising version of this mode, namely that for maximum angular resolution.

A routine real-time procedure, CAP 631, was carried out to update the on-board ephemeris coefficients.

The planned observations for the coming week are as follows.

------------------------------------------
M31 monitoring (5 obs) HRC-I Jul 16 R Aquarii ACIS-S Radiation Belt Jul 17 NGC 4486 ACIS-S/HETG GX 9+1 ACIS-S/HETG GX 5-1 ACIS-S/HETG Jul 18 star TW Hya ACIS-S/HETG X0918-54 ACIS-S/HETG Jul 19 LMC SNR N132D ACIS-S/HETG Radiation Belt LMC SNR N132D ACIS-S/HETG Jul 20 HD 45348 ACIS-S/HETG Jul 21 Radiation Belt Jul 22 Seyfert, PG 1404+226 ACIS-S/HETG Composite SNR KES 73 ACIS-S Jul 23

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports